Status Report

SOHO Daily Meeting Minutes for Friday, 28 December 2001

By SpaceRef Editor
December 28, 2001
Filed under , ,

* Holiday contact information – Each instrument team should email
to the SOCs a special contact list with vacation dates and phone
numbers where people can be reached while out of town.

* It may be impossible to access MDI data on 30-31 December since they are
re-arranging the computer room.

* DSS-16 is in maintenance and won’t be up until week 2 of 2002.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, SWAN, Nominal HGA Table,
New tracking star (Xmas star)

Planned Activities: VIRGO, CELIAS, RSL
New control star, mag = 2.8 (Xmas star)
New tracking star, mag = 7.1

Upcoming Operations:
1/2/02 CDMU/ACU Memory Dumps

Ground Anomalies:
361/2008 Lost uplink lock on spacecraft RCVR 2, cause
unknown. Lost command ability for 4 minutes.
DR #C100760.

12/27 DOY 361 – 12/28 DOY 362

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Active region variability study continues.
Supporting JOP151.

UVCS: Nominal. JOP151 Active Region Streamer Study on AR 9742
and synoptics.

LASCO: Nominal. C2/C3 synoptics supporting JOP151. For the next month,
color sequences will be taken daily in C3 for flat fielding purposes.

EIT: Nominal. 195 Angstrom CME watch, synoptics, and JOP151.

MDI: Nominal. Full disk campaign of magnetograms and dopplergrams.
Supporting JOP151.

TRACE: Nominal. Observing AR 9742 today.

SpaceRef staff editor.